# Raleigh, NC 27606, USA

### v1

@@ 1 Introduction
As renewable energy sources become more prevalent and grid infrastructure continues to age, the need for cost-effective planning and operational decisions is becoming increasingly critical. Transmission Expansion Planning (TEP) is used to make long-term upgrades [1], but its deployment is often limited by capital, regulatory, and siting constraints [2]. As a complementary strategy, topology control through transmission line switching has emerged as a systematic approach to improving the operational efficiency of existing infrastructure, as a fixed topology is rarely optimal across all time periods or operating conditions [3]. This strategy is mathematically formulated as the Optimal Transmission Switching (OTS) problem, which determines an efficient network configuration by selectively switching transmission lines on or off [4]. Its benefits are well documented and include improved congestion management and controllability, along with reduced operating costs [5, 6, 7].
OTS extends Optimal Power Flow (OPF)—which minimizes generation costs subject to physical and operational constraints—by incorporating binary line-status variables that modify the network topology. The highest-fidelity OPF, Alternating Current OPF (ACOPF), is computationally challenging and offers only local optimality guarantees, as it is non-convex [8]. Adding the discrete decisions of OTS further exacerbates this intractability. Thus, various studies adopt the linearized Direct Current OPF (DCOPF) model (e.g., [9, 10, 11, 12]), which simplifies the physics yet retains relevance to practice. DCOPF-based OTS (DC-OTS) introduces binary variables for reconfiguring the network topology, thereby transforming DCOPF from a linear program (LP) into a mixed-integer linear program (MILP). Although computationally more manageable than its ACOPF counterpart, DC-OTS remains NP-hard [13]. The problem is typically modeled with “big-MM” formulations (e.g., [9],[11],[14]), most notably the angle-based model of Fisher et al. [4]. This formulation is challenging to scale, largely due to its disjunctive constraints. We construct an extended formulation for an angle-based relaxation of this model and, using projection, derive valid inequalities (VIs) that describe its convex hull.
@@ 1.1 Model and related work
Big-$M$ parameters ($M_{ij}$ in (1d)) impose an upper bound on the voltage angle difference between buses across inactive transmission lines. These constants must be chosen with care: They should be large enough to preserve all integer-feasible solutions (since optimal line statuses are not known a priori), yet sufficiently tight to avoid overly conservative values that can lead to weak relaxations, increasing the computational difficulty of solving DC-OTS exactly [9].
To avoid large big-$M$ values, previous works adopt fixed angle-difference bounds—such as $2\pi$ or $1.2$ radians (e.g., [9, 15, 16, 17])—partly motivated by operational considerations. However, in DCOPF, angles are treated as real-valued variables (“non-periodic”), and thus imposing fixed bounds can at times exclude feasible solutions [18]. Furthermore, these bounds are usually defined only for adjacent buses and provide no guidance for the $\Theta(|\mathcal{B}|^{2})$ non-adjacent pairs in the network. Even with fixed bounds, the computational complexity arising from the combinatorial nature of the solution space remains. This persistent complexity, coupled with the limited scalability of traditional algorithms, has led to an increased reliance on heuristic and data-driven approaches (e.g., [10, 19, 20, 21, 22]).
In power systems with flexible network topologies, it is often assumed that the NP-hard longest path problem (LPP) needs to be solved to tighten the angle-difference bounds [18, 23]. To avoid solving the LPP exactly, Aguilar-Moreno et al. [24] simplify the network and use an LPP relaxation to set big-$M$ constants; however, this approach can further overestimate effective transfer paths, resulting in big-$M$ values that remain excessively loose.
@@ 2 Preliminaries
One workaround is to use a trivial bound for any flexible topology by summing the weights of all lines in the network [28]. Although easy to compute, this bound ignores network structure and is typically very loose. To obtain tighter bounds, Binato et al. [23] propose total enumeration, in which all potential paths connecting two buses are evaluated and the longest is selected, guaranteeing validity regardless of switching decisions. However, this approach is equivalent to solving the LPP, which is NP-hard [29]. Moreover, Fattahi et al. [18] show that, unless $P=NP$, even approximating the maximum feasible angle difference is intractable. Beyond these computational challenges, LPP-based bounds often yield big-$M$ values that substantially overstate realistic angle differences.

### v2

# Facet-Defining Inequalities for the Angle-Based DC Optimal Transmission Switching Formulation

@@ 1 Introduction
As renewable energy sources become more prevalent and grid infrastructure continues to age, the need for cost-effective planning and operational decisions is becoming increasingly critical. Transmission Expansion Planning (TEP) is used to make long-term upgrades [1], but its deployment is often limited by capital, regulatory, and siting constraints [2]. A complementary strategy is topology control via transmission line switching, which has emerged as a systematic approach for improving the operational efficiency of existing infrastructure. This is motivated by the fact that a fixed topology is rarely optimal across all time periods or operating conditions [3]. This strategy is mathematically formulated as the Optimal Transmission Switching (OTS) problem, which determines an efficient network configuration by selectively switching transmission lines on or off [4]. Its benefits are well documented and include improved congestion management and controllability, along with reduced operating costs [5, 6, 7].
OTS extends Optimal Power Flow (OPF)—which minimizes generation costs subject to physical and operational constraints—by incorporating binary line-status variables that modify the network topology. The highest-fidelity OPF, Alternating Current OPF (ACOPF), is computationally challenging and offers only local optimality guarantees, as it is non-convex [8]. Introducing discrete line-switching decisions further exacerbates this intractability. Thus, various studies adopt the linearized Direct Current OPF (DCOPF) model (e.g., [9, 10, 11, 12]), which simplifies the physics yet retains relevance to practice. DCOPF-based OTS (DC-OTS) introduces binary variables for reconfiguring the network topology, thereby transforming DCOPF from a linear program (LP) into a mixed-integer linear program (MILP). Although computationally more manageable than its ACOPF counterpart, DC-OTS remains NP-hard [13]. The problem is typically modeled with “big-MM” formulations (e.g., [9],[11],[14]), most notably the angle-based model of Fisher et al. [4]. This formulation is challenging to scale, largely due to its disjunctive constraints. We construct an extended formulation for an angle-based relaxation of this model and, using projection, derive valid inequalities (VIs) that describe its convex hull.
@@ 1.1 Model and related work
Big-$M$ parameters ($M_{ij}$ in (1d)) impose an upper bound on the voltage angle difference between buses across inactive transmission lines. These constants must be large enough to preserve all integer-feasible solutions (since the optimal line statuses are not known a priori), yet sufficiently tight to avoid weak LP relaxations that can substantially increase the computational burden of solving DC-OTS exactly [9].
To avoid large big-$M$ values, previous works adopt fixed angle-difference bounds—such as $2\pi$ or $1.2$ radians (e.g., [9, 15, 16, 17])—partly motivated by operational considerations. However, DCOPF models phase angles as real-valued (“non-periodic”) variables, and thus imposing fixed bounds can at times exclude feasible solutions [18]. In addition, these bounds are typically specified only for adjacent buses and provide no guidance for the $\Theta(|\mathcal{B}|^{2})$ non-adjacent pairs in the network. Even with fixed bounds, the computational complexity arising from the combinatorial nature of the solution space remains. This persistent complexity, coupled with the limited scalability of traditional algorithms, has led to an increased reliance on heuristic and data-driven approaches (e.g., [10, 19, 20, 21, 22]).
In power systems with flexible network topologies, it is often assumed that the NP-hard longest path problem (LPP) needs to be solved to tighten the angle-difference bounds [18, 23]. To avoid solving the LPP exactly, Aguilar-Moreno et al. [24] simplify the network and solve a relaxed problem to derive big-$M$ constants. However, the relaxation can further overestimate effective transfer paths, yielding big-$M$ values that remain excessively loose.
@@ 2 Preliminaries
One workaround is to use a trivial bound for any flexible topology by summing the weights of all lines in the network [28]. Although easy to compute, this bound ignores network structure and is typically very loose. To obtain tighter bounds, Binato et al. [23] propose total enumeration of all simple paths connecting two buses to identify the longest one, ensuring validity under any switching decision. However, this approach is equivalent to solving the LPP, which is NP-hard [29]. Moreover, Fattahi et al. [18] show that, unless $P=NP$, even approximating the maximum feasible angle difference is intractable. Beyond these computational challenges, LPP-based bounds often yield big-$M$ values that substantially overstate realistic angle differences.